OpenStudy (anonymous):

Jake is comparing the prices of two laundry service companies. Company A charges $20 per load and an additional $8 as service charges. Company B charges $24 per load and an additional $4 as service charges. Part A: Write equations to represent Company A and Company B's total laundry charges for a certain number of loads. Define the variable used in the equations. (4 points) Part B: Which company would charge less for 5 loads of laundry? Justify your answer. (3 points) Part C: How much money is saved by using the services of Company A instead of Company B to clean 9 loads of laundry?

OpenStudy (anonymous):

Part A: Company A: 20x+8 Company B: 24x+4 X is the number of loads Part B: Company A: 20(5)+8 = $108 Company B: 24(5)+4= $124 So, company A would charge less. Part C: Company A: 20(9)+8= $188 Company B: 24(9)+4= $220 $32 is the amount saved